# Language Models as Critical Thinking Tools: A Case Study of Philosophers

### v1

@@ Abstract
Current work in language models (LMs) helps us speed up or even skip thinking by accelerating and automating cognitive work. But can LMs help us with critical thinking – thinking in deeper, more reflective ways which challenge assumptions, clarify ideas, and engineer new concepts? We treat philosophy as a case study in critical thinking, and interview 21 professional philosophers about how they engage in critical thinking and on their experiences with LMs. We find that philosophers do not find LMs to be useful because they lack a sense of selfhood (memory, beliefs, consistency) and initiative (curiosity, proactivity). We propose the selfhood-initiative model for critical thinking tools to characterize this gap. Using the model, we formulate three roles LMs could play as critical thinking tools: the Interlocutor, the Monitor, and the Respondent. We hope that our work inspires LM researchers to further develop LMs as critical thinking tools and philosophers and other ‘critical thinkers’ to imagine intellectually substantive uses of LMs.
@@ 1 Introduction
Language Models (LMs) have recently alleviated a whole host of our intellectual inconveniences. They can help us do things we would have begrudgingly done by ourselves otherwise: write code (Chen et al., 2021; Rozière et al., 2024), generate emails  (Goodman et al., 2022), and translate text  (Costa-jussà et al., 2022). In sparking ideas by generating stories (Schwitzgebel et al., 2023) and concept designs (Ma et al., 2023), LMs offer shortcuts to gaining new thoughts. They also help us put our thinking into words by revising (Mysore et al., 2023) and giving feedback (Liang et al., 2023) on our writing. In all these cases, LMs help us speed up and circumvent the inconveniences of thinking ourselves.
In many contexts, however, the “inconvenience” of thinking is not a temporary problem to be alleviated, but a deep puzzle to be reflected upon. Many people are invested in specific areas of intellectual inquiry – e.g., historians, scientists, philosophers – and more generally, in reflection and engagement with the world – e.g., as informed political citizens, critical consumers of media, and moral actors. They are interested in identifying and challenging assumptions, clarifying muddled ideas, and engineering new and useful ways to think. Core to this sort of inquiry is critical thinking – “the propensity and skill to engage in an activity with reflective skepticism” (McPeck, 2016). Can LMs serve as tools for critical thinking – helping us think more deeply and in more complex ways, rather than faster or not at all? What if – like John – we claim all the rights to think?
To investigate how LMs can serve as critical thinking tools, we use philosophers as a case study – philosophers being people who are in the business of thinking critically about a wide range of concepts and ideas. We interview 21 professional philosophers to understand their thinking processes, collect their experiences with and views on current LMs, and brainstorm the roles LMs could play as critical thinking tools in philosophy (§3). We find that current philosophers do not think LMs are good critical thinking tools (§4) for two primary reasons: LMs are too neutral, detached, and nonjudgmental (§4.2); and LMs are too servile, passive, and incurious (§4.3). We propose the selfhood-initiative model for critical thinking tools, which explains why philosophers find conversations with other philosophers and reading philosophical texts to be more helpful for their work than current LMs (§5.1). Using the model, we describe three roles LMs could play as critical thinking tools: the Interlocutor, the Monitor, and the Respondent (§5.2). Finally, we outline how these LMs could inform metaphilosophical questions and shape the discipline of philosophy (§6.1), and discuss challenges in building LMs (§6.2) and interfaces (§6.3) for critical thinking.
@@ 2.1 Language Models as Thinking Tools
A large and growing literature investigates how LMs can serve as thinking tools for humans engaged in intellectual work. This research tends to concern how LMs can serve two intellectual functions: idea stimulation and idea expression.
LMs can provide stimulus for ideas – information and formulations which provoke and guide creative processes. In these roles, LMs expand the set of ideas a user has available, which can become inspiration for thinking. These LMs may continue to provide stimulus while the user is thinking. For instance, LMs can produce creative analogies (Bhavya et al., 2023), metaphors (Chakrabarty et al., 2021), ‘sparks’ (Gero et al., 2021) to stimulate serious thinking. Existing pools of ideas might be expanded with LMs (Fede et al., 2022). Creative writers may use LMs to propose plots, characters, and entire stories (Yuan et al., 2022; Mirowski et al., 2022; Schmitt & Buschek, 2021; Calderwood et al., 2020; Chakrabarty et al., 2023); designers may use LMs to generate concept designs (Ma et al., 2023); scientists may use LMs to find literature (Morris, 2023) and (re)formulate problems in novel ways (Gu, 2023).
On the other hand, LMs can also aid the expression of ideas (in language). It is assumed that users have an idea in mind and wish to use LMs to find the formulation which best represents it. LMs can help ‘translate ideas into words’ by making suggested revisions (Du et al., 2022; Mysore et al., 2023; Zhao, 2022; Shu et al., 2023; Lin et al., 2024) and helping the user clarify their writing goals (Arnold et al., 2021; Kim et al., 2024). Writing feedback given by LMs may be more useful (Liang et al., 2023), more motivating (Meyer et al., 2024), and more engaging (Tanwar et al., 2024) than feedback given by other humans.
@@ 2.2 Language Models as Critical Thinking Tools
However, one part of the thinking process is clearly missing. One does not simply go from the stimulus for an idea to figuring out how to express the idea: one needs to do the actual critical thinking, involving reflection upon ideas, judgment, and conceptual engineering. LMs can help provide the seeds for our ideas when we don’t have any (i.e., stimulus) and help us formulate them once we’ve got them (i.e., expression), but how can they help us with questioning, reorienting, analyzing, and developing ideas (i.e., critical thinking)?
@@
There are many different definitions of critical thinking: “the propensity and skill to engage in an activity with reflective skepticism” (McPeck, 2016), “reasonable, reflective thinking that is focused on deciding what to believe or do” (Ennis, 1993), and “the development and evaluation of arguments” (Facione, 1984). Critical thinking requires many dispositions, such as seeking clear statements of questions, looking for alternatives, and being open-minded (Ennis, 1987). Critical thinking is what makes many areas of intellectual inquiry – such as history, science, and philosophy – difficult. For instance, on different accounts, history requires interpreting the past with alternative (nonlinear, long-range) temporalities (Braudel, 2023), taking into account the ways in which power structures shape historical record (Foucault, 1995; Trouillot, 1995), and identifying and manipulating narrative structures (White, 1975; Gaddis, 2004). Science requires advances not only in empirical work, but also in the underlying paradigms of research (Kuhn & Hawkins, 1963) and navigation of a variety of social, technical, and material factors (Latour, 1989).
@@
A limited body of work has explored how LMs might contribute towards human critical thinking. Rodman (2023) suggests that political theorists might use LMs to metacognitively reflect upon their own creative processes and judgments. Cai et al. (2024) consider how currently “sycophantic”, “servile”, and “lobotimized” LMs can be used in more critical ways by challenging users’ pre-existing ideas and constructively using antagonistic interactions to develop their thinking. Park & Kulkarni (2024) show how LMs as “thinking assistants” can facilitate human self-reflection by asking, instead of answering, questions.
2.3 Critical Thinking as Philosophy
@@
In this paper, we focus on philosophy as a case study for critical thinking. Philosophy is concerned with critical, systematic, and reflective examination of the world. This includes understanding the basic structure of life and the world – what does it mean to exist (Aristotle, 1908; Heidegger, 1962; Sartre, 1993), live (Aurelius, 2006), and die (Kierkegaard, 1983; Nietzsche, 1961)?; what does it mean to know something (Plato, 2008b; Kant, 2003; Husserl, 1977) and what are the limits of scientific knowledge  (Popper, 2002; Chalmers, 2013)?; on what moral bases should we act (Aristotle, 2004; Spinoza, 2003), and is it even possible to determine ‘objective’ answers to moral questions (Hume, 2003; Harman & Thomson, 1996)? Core to philosophy is “the endeavour to know how and to what extent it might be possible to think differently, instead of legitimating what is already known” (Foucault & Hurley, 1990). In thinking about how to think, philosophy is not only about suspicion toward the meanings and functions of various phenomena, but also recovery of new significances, meanings, and coherence (Ricoeur & Thompson, 1981).
@@
Contrary to the image that philosophy is “done in the armchair”, isolated and impractical, philosophy has always been intertwined with other lines of inquiry. Plato engaged extensively with advanced mathematics; Aristotle contributed to early physics; Hume leaned on psychology. Philosophy has asked and continues to ask urgent, relevant questions: how are we to understand “fairness” and “justice” in algorithmic discrimination (Hu, 2023), legal punishment (Alexander, 1922), and the distribution of resources (Rawls, 1971); quantum mechanics in physics (Carnap, 1966); the relationship between consciousness and the brain in neuroscience (Chalmers, 2013)? Indeed, researchers in every area of intellectual inquiry confront philosophical questions in their work: they might ask if a model or concept is “really real” and how they know so, or aim to formulate normative desiderata for theories.
@@ 3 Methods
The first author conducted interviews with 21 professional philosophers at 14 philosophy departments at doctoral universities in the United States. We contacted and selected philosophers for high diversity across area of interest (e.g., ethics, political philosophy, philosophy of science). Interviews took place online and lasted between 30 to 60 minutes, depending on interviewee availability. Interviewees were asked how they philosophize (e.g., where ideas come from, how ideas are developed, what resources are needed) and their views on LMs (e.g., can LMs ‘do’ philosophy, how might they be useful for philosophizing). These questions followed a loose script (see §A.1), although we asked novel follow-up questions to pursue interesting lines of inquiry raised by the interviewees’ responses. In cases where interviewees had very little or no prior exposure to LMs, they interacted live with the GPT-4 model on a philosophical topic of their choosing. We received IRB approval from our university to conduct the interviews; all interviewees confirmed their consent to participate in the study, and for their responses to inform the development of this article. We qualitatively analyzed interview recordings and transcripts. Using an inductive approach (Thomas, 2006) and open coding (Charmaz, 2006), we identified common themes and positions (yielding §4 and §5). We refer to interviewees with a unique identifier, e.g., (P1, P2, P3).
@@ 4 Language Models Are Not Good Critical Thinking Tools (So Far)
Many of the interviewed philosophers find LMs to be relevant and interesting, and some find them to have limited uses such as for undergraduate instruction (P1, P13, P20) or becoming acquainted with a topic (P5, P11, P12). However, none of the philosophers were convinced that current LMs can reliably and conveniently assist them in the intellectually substantive ways which require critical thinking. Philosophers described current LMs as “boring” (P2), “anodyne” (P4), “bland” (P9), and “cowardly” (P13). There are two broad reasons for this. First, current LMs tend to be highly neutral, detached, and non-judgmental, often commenting on ideas in abstract and decontextualized ways (§4.2). Second, current LMs tend to be servile, passive, and incurious, which is unhelpful when the user does not yet have a clear vision of what they want to accomplish; this restricts the variety of intellectual interactions possible S4.3).
@@ 4.1 How do philosophers philosophize?
Where do philosophical ideas come from? Philosophers report that their ideas usually come from observing puzzles and tensions in the world, in which some aspect feels bothersome (P5, P12, P20), incomplete (P10, P14), in need of clarity (P1, P13), or outright incorrect (P3). Philosophers encounter these puzzles and tensions most commonly in open conversation with others (P1, P2, P5, P9, P19) and while reading texts – books, papers, and monographs making explicitly philosophical arguments or touching upon philosophical themes (P4, P7, P10, P12, P13, P20). These puzzles may have an intellectual or logical character: terms might not be sufficiently disambiguated, inferences may not be valid, and propositions may entail absurd conclusions (P8, P11). However, for many, these tensions are identified and drawn out by ethical motivations (P1, P8, P16, P12). Tensions might arise not primarily because a proposition is incoherent, but rather because it appears ethically problematic. The now-famous trolley problem dilemma was used to probe the differences between doing and allowing harm, with applications to bioethics, particularly abortion (Foot, 1967). Several philosophers describe being inspired by texts communicating empirical work, seeking to provide explanations for empirical observations (P1, P2, P16, P18) as well as subjecting the practices and products of the empirical sciences to critical inquiry (P2, P7, P12, P13, P18).
What do philosophers want out of their ideas? Once philosophers identify puzzles from conversations and texts, they aim to develop ideas which make progress on these puzzles. Progress is conceived of in many ways: “understand[ing] some part of the world better” (P3), working through new ways to think about problems (P17), and better understanding the current ways we think – for instance, by making implicit assumptions explicit and recognizing the implications of propositions (P7). Some philosophers describe a developed philosophical idea as a “picture” (P9, P10) which organizes subideas in a systematic way, allowing one to clearly see the main point(s). This often requires “conceptual engineering” (P6): challenging, disassembling, and rebuilding the ways in which we think.
@@
The role of conversation in philosophical development. Conversations with fellow philosophers are central to evaluating the coherence of ideas (P21), raising connections to other ideas and problems (P5), and collecting criticism, objections, and feedback (P3, P10). Conversations may force philosophers to explain and justify ideas they may have taken for granted (P1). Conversation helps philosophers gain confidence that their ideas are good intellectual contributions (P2, P21). Philosophers even simulate conversations in their head, taking on various positions for and against their ideas (P1, P12). Good philosophical conversation requires several conditions. The interlocutor should be charitable – genuinely listening to and working through ideas (P1, P12), and trusting (P6, P14) – but also willing to boldly push ideas forward (P3) and take intellectual risks (P18). Conversations may not be directed towards any clear goal; interlocutors must be able to “riff off each other” (P8) and be willing to operate without a preset agenda (P3, P4). This requires interlocutors to be curious about addressing problems (P21); it should be a collaborative enterprise, rather than a combative debate (P3, P7).

### v2

@@ Abstract
Current work in language models (LMs) helps us speed up or even skip thinking by accelerating and automating cognitive work. But can LMs help us with critical thinking — thinking in deeper, more reflective ways which challenge assumptions, clarify ideas, and engineer new concepts? We treat philosophy as a case study in critical thinking, and interview 21 professional philosophers about how they engage in critical thinking and on their experiences with LMs. We find that philosophers do not find LMs to be useful because they lack a sense of selfhood (memory, beliefs, consistency) and initiative (curiosity, proactivity). We propose the selfhood-initiative model for critical thinking tools to characterize this gap. Using the model, we formulate three roles LMs could play as critical thinking tools: the Interlocutor, the Monitor, and the Respondent. We hope that our work inspires LM researchers to further develop LMs as critical thinking tools and philosophers, and other ‘critical thinkers’ to imagine intellectually substantive uses of LMs.
[…]
[leftline=true, linecolor=black, topline=false, bottomline=false, rightline=false]leftbarquote
@@ 1 Introduction
Language Models (LMs) have recently alleviated a whole host of our intellectual inconveniences. They can help us do things we would have begrudgingly done by ourselves otherwise: write code (Chen et al., 2021; Rozière et al., 2023), generate emails  (Goodman et al., 2022), and translate text  (Costa-jussà et al., 2022). In sparking ideas by generating stories (Schwitzgebel et al., 2023) and concept designs (Cai et al., 2023), LMs offer shortcuts to gaining new thoughts. They also help us put our thinking into words by revising (Mysore et al., 2023) and giving feedback (Liang et al., 2024) on our writing. In all these cases, LMs help us speed up and circumvent the inconveniences of thinking ourselves.
In many contexts, however, the “inconvenience” of thinking is not a temporary problem to be alleviated, but a deep puzzle to be reflected upon. Many people are invested in specific areas of intellectual inquiry — e.g., historians, scientists, philosophers — and more generally, in reflection and engagement with the world — e.g., as informed political citizens, critical information consumers, and moral actors. They are interested in identifying and challenging assumptions, clarifying muddled ideas, and engineering new and useful ways to think. Core to this sort of inquiry is critical thinking — “the propensity and skill to engage in an activity with reflective skepticism” (McPeck, 2016). Can LMs serve as tools for critical thinking — helping us think more deeply and in more complex ways, rather than faster or not at all? What if — like John — we claim all the rights to think (Buçinca et al., 2021)?
To investigate how LMs can serve as critical thinking tools, we use philosophers as a case study — philosophers being people who are in the business of thinking critically about a wide range of concepts and ideas. We interview 21 professional philosophers to understand their thinking processes, collect their experiences with and views on current LMs, and brainstorm the roles LMs could play as critical thinking tools in philosophy (§3). We find that current philosophers do not think LMs are good critical thinking tools (§4) for two primary reasons: LMs are too neutral, detached, and nonjudgmental (§4.2); and LMs are too servile, passive, and incurious (§4.3). We propose the selfhood-initiative model for critical thinking tools, which explains why philosophers find conversations with other philosophers and reading philosophical texts to be more helpful for their work than current LMs (§5.1). Using the model, we describe three roles LMs could play as critical thinking tools: the Interlocutor, the Monitor, and the Respondent (§5.2). Finally, we outline how these LMs could inform metaphilosophical questions and shape the discipline of philosophy (§6.3), and discuss challenges in building LMs (§6.1) and interfaces (§6.2) for critical thinking.
@@ 2.1 Language Models as Thinking Tools
A large and growing literature investigates how LMs can serve as thinking tools for humans engaged in intellectual work. This research tends to concern how LMs can serve two intellectual functions: idea stimulation (roughly, “divergent thinking”) and idea refinement (roughly, “convergent thinking”) (Banathy, 1996; Design Council, 2019).
LMs can provide stimulus for ideas — information and (re)formulations which provoke and guide creative processes. Generally, LMs can expand idea sets (Fede et al., 2022), produce creative analogies (Bhavya et al., 2023) and metaphors (Chakrabarty et al., 2021); discover concepts (Lam et al., 2024), and facilitate group ideation (Rayan et al., 2024; Shaer et al., 2024). LMs to open-endedly propose plots, characters, and entire stories for creative writers (Calderwood et al., 2020; Schmitt & Buschek, 2021; Yuan et al., 2022; Mirowski et al., 2023; Chung et al., 2022; Chakrabarty et al., 2023); but also provide inspiration in more constrained creativity tasks, such as science writing (Gero et al., 2021; Kim et al., 2023). Although fraught with pitfalls (Messeri & Crockett, 2024), scientists can use LMs to find and synthesize literature (Van Dinter et al., 2021; Wagner et al., 2022; Fok et al., 2023; Khraisha et al., 2024) and iterate through research inquiry (Wang et al., 2023a; Morris, 2023; Liu et al., 2024). Designers can use LMs to generate and develop concept designs  (Cai et al., 2023; Liu et al., 2023; Ma et al., 2023; Chong et al., 2024; Ma et al., 2024; Chen et al., 2024).
On the other hand, LMs can also aid the refinement of ideas – selecting from and improving upon an established existing pool of ideas. LMs can help writers by making suggested revisions (Du et al., 2022; Zhao, 2022; Mysore et al., 2023; Shu et al., 2023) and clarifying writing goals (Arnold et al., 2021; Kim et al., 2024). For scientists, LMs can facilitate revision of scientific writing (Liang et al., 2024; Radensky et al., 2024); for designers, LMs can provide feedback on (Duan et al., 2024) and annotate (Lu et al., 2024) UIs. In teaching settings, writing feedback given by LMs may be more motivating (Meyer et al., 2024) and engaging (Tanwar et al., 2024) than feedback given by other humans. Besides reviewing ideas, LMs can also curate them — for instance, by summarizing writing (Fabbri et al., 2021; Dang et al., 2022) and identifying important ideas (Lin et al., 2024).
@@ 2.2 Language Models as Critical Thinking Tools
However, one part of the thinking process is clearly missing. One does not simply go from the stimulus for ideas to figuring out how to refine them: one needs to do the actual critical thinking, involving reflection upon ideas, judgment, and conceptual engineering. LMs can help provide the seeds for our ideas when we don’t have any (i.e., stimulus) and help us work through them once we’ve got them (i.e., refinement), but how can they help us with questioning, reorienting, analyzing, and developing ideas (i.e., critical thinking)?
@@
There are many different definitions of critical thinking: “the propensity and skill to engage in an activity with reflective skepticism” (McPeck, 2016), “reasonable, reflective thinking that is focused on deciding what to believe or do” (Ennis, 1993), and “the development and evaluation of arguments” (Facione, 1984), among many others. Critical thinking requires many dispositions, such as seeking clear statements of questions, looking for alternatives, and being open-minded (Ennis, 1987). Critical thinking is what makes many areas of intellectual inquiry — such as history, science, and philosophy — difficult. In these areas, people must produce and work with observations that are incomplete and open to a multiplicity of framings to pursue problems with often unclear definitions of progress — a landscape which demands critical thinking. For instance, on different accounts, history requires interpreting the past with alternative (nonlinear, long-range) temporalities (Braudel, 2023), taking into account the ways in which power structures shape historical record and memory (Foucault, 1969a; Trouillot, 1995), and identifying and manipulating narrative structures (White, 1973; Gaddis, 2004). Science requires advances not only in empirical work, but also reflection upon underlying paradigms of research (Kuhn & Hawkins, 1963), epistemology (Harding, 2013), and the social and material factors that constitute scientific knowledge (Latour, 1989).
Researchers across a variety of fields have developed a rich tapestry of approaches and tools to support critical thinking and related acts. Educators develop teaching strategies to promote critical thinking (McPeck, 1990; Pithers & Soden, 2000) such as teaching and interlinking a variety of perspectives on a subject in an integrative manner (Enciso et al., 2017) and encouraging students’ intellectual independence in finding answers to their questions (Langer, 1997; Raths et al., 1966). Psychologists and cognitive scientists seek to understand how cognitive mechanisms and biases inform how humans (should) develop critical thinking (Carey, 1986; Reif, 2008), emphasizing the cultivation of basic metacognitive “building blocks” of critical thinking (Pasquinelli et al., 2021) and teaching for “practical theory” (Gelder, 2005). Meanwhile, human-computer interaction (HCI) researchers explore how interactions with computer applications can facilitate critical thinking: designers can provoke experiences of discomfort (Benford et al., 2012; Halbert & Nathan, 2015); emphase understanding over rote expression in social contexts (Kriplean et al., 2012; Sun et al., 2017); and build small “nudges” (Liao & Wang, 2022) into interfaces which “prime” (Yamamoto & Yamamoto, 2018) users towards reflective critical thinking (Bentvelzen et al., 2022); among many others. Many of these themes will be revisited in our discussion of design proposals for LMs as critical thinking tools (§5.2).
@@
A growing body of work has explored how LMs might contribute towards critical thinking. LM-based news and media can positively affect users’ willingness to think through opposing or novel viewpoints, which can be applied to combat polarization and extremism (Tanprasert et al., 2024; Zarouali et al., 2021; Shin, 2022; Wang & Tanes-Ehle, 2022; Blasiak et al., 2021). Cai et al. (2024) consider how currently “sycophantic”, “servile”, and “lobotimized” LMs can be used in more critical ways by challenging users’ pre-existing ideas and constructively using antagonistic interactions to develop their thinking. Danry et al. (2023); Ma et al. (2023); Park & Kulkarni (2023) show how LMs can facilitate human self-reflection and improve human reasoning by asking questions instead of only answering them (as in the typical LM interaction paradigm). Xu et al. (2024) encourage critical thinking by building LM interactions using structured templates (over free-form chat). In more targeted contexts, LMs can be used to help scientific researchers critically think about their impact statements (Mukherjee et al., 2023), and to help political theorists to metacognitively reflect upon their own creative processes and judgments (Rodman, 2023).
2.3 Philosophy as Critical Thinking, Critical Thinking as Philosophy
@@
In this paper, we focus on philosophy as a case study for critical thinking. Philosophy is concerned with critical, systematic, and reflective examination of the world. This includes understanding the basic structure of life and the world — what does it mean to exist (Aristotle, 350 BCE; Heidegger, 1927; Sartre, 1943), live (Aurelius, 180 AD), and die (Kierkegaard, 1983; Nietzsche, 1892)?; what does it mean to know something (Plato, 369BCE; Kant, 1781; Husserl, 1931) and what are the limits of scientific knowledge  (Popper, 2002; Chalmers, 2013)?; on what moral bases should we act (Aristotle, 350BCE; Spinoza, 1677), and is it even possible to determine ‘objective’ answers to moral questions (Hume, 1739; Harman & Thomson, 1996; Foucault, 1976)? Core to philosophy is “the endeavour to know how and to what extent it might be possible to think differently, instead of legitimating what is already known” (Foucault, 1976). Philosophy is for intellectual creation and engineering: Deleuze & Guattari (1991) wrote that “So long as there is a place for creating concepts, the operation that undertakes this will always be called philosophy.” In thinking about how to think, philosophy is not only about suspicion toward the meanings and functions of phenomena, but also recovery of new significances and coherence (Ricoeur, 1981).
@@
Contrary to the image that philosophy is “done in the armchair”, isolated and impractical, philosophy has always been intertwined with other lines of inquiry. Plato engaged extensively with advanced mathematics; Aristotle contributed to early physics; Hume leaned on psychology. Philosophy has asked and continues to ask urgent, relevant questions: for instance, how are we to understand the strangeness of quantum mechanics in physics (Carnap, 1966); the relationship between consciousness (mind) and the brain (matter) (Chalmers, 2013); and “fairness” and “justice” in contexts like algorithmic discrimination (Hu, 2023), legal punishment (Alexander, 1922), and the distribution of resources (Rawls, 1971)? Indeed, researchers in every area of intellectual inquiry confront philosophical questions in their work: they might ask if a model or concept is “really real” (and how they know so), what the “nature” of their object of study is, aim to formulate normative desiderata for their theories, and so on. Therefore, we study philosophers’ views and practices in this paper both because philosophers engage extensively in critical thinking and because many questions which require critical thinking asked by non-philosophers often have a philosophical flavor.
@@ 3 Methods
The first author conducted interviews with 21 professional philosophers at 14 philosophy departments at doctoral universities in the United States. We contacted and selected philosophers for high diversity across area of interest (e.g., ethics, political philosophy, philosophy of science). Interviews took place online and lasted between 30 to 60 minutes, depending on interviewee availability. Interviewees were asked how they philosophize (e.g., where ideas come from, how ideas are developed, what resources are needed) and their views on LMs (e.g., can LMs ‘do’ philosophy, how might they be useful for philosophizing). These questions followed a loose script (see §B), although we asked novel follow-up questions to pursue interesting lines of inquiry raised by the interviewees’ responses. In cases where interviewees had very little or no prior exposure to LMs, they interacted live with the GPT-4 model on a philosophical topic of their choosing. We received IRB approval from our university to conduct the interviews; all interviewees confirmed their consent to participate in the study, and for their responses to inform the development of this paper. We qualitatively analyzed interview recordings and transcripts. Using an inductive approach (Thomas, 2006) and open coding (Charmaz, 2006), we identified common themes and positions (yielding §4 and §5). We refer to interviewees with a unique identifier, e.g., (P1, P2, P3) (see §A).
@@ 4 Language Models Are Not Good Critical Thinking Tools (So Far)
Many of the interviewed philosophers find LMs to be relevant and interesting, and some find them to have limited uses such as for undergraduate instruction (P1, P13, P20) or becoming acquainted with a topic (P5, P11, P12). However, none of the philosophers were convinced that current LMs can reliably and conveniently assist them in the intellectually substantive ways which require critical thinking. Philosophers described current LMs as “boring” (P2), “anodyne” (P4), “bland” (P9), and “cowardly” (P13). We discovered two broad reasons for this. First, current LMs tend to be highly neutral, detached, and non-judgmental, often commenting on ideas in abstract and decontextualized ways (§4.2). Second, current LMs tend to be servile, passive, and incurious, which is unhelpful when the user does not yet have a clear vision of what they want to accomplish, restricting the variety of intellectual interactions possible  S4.3).
@@ 4.1 How do philosophers philosophize?
Where do philosophical ideas come from? Philosophers report that their ideas usually come from observing puzzles and tensions in the world, in which some aspect feels bothersome (P5, P12, P20), incomplete (P10, P14), in need of clarity (P1, P13), or outright incorrect (P3). Philosophers encounter these puzzles and tensions most commonly in open conversation with others (P1, P2, P5, P9, P19) and while reading texts — books, papers, and monographs making explicitly philosophical arguments or touching upon philosophical themes (P4, P7, P10, P12, P13, P20). These puzzles may have an intellectual or logical character: terms might not be sufficiently disambiguated, inferences may not be valid, and propositions may entail absurd conclusions (P8, P11). However, for many, these tensions are identified and drawn out by ethical motivations (P1, P8, P16, P12). Tensions might arise not primarily because a proposition is incoherent, but rather because it appears ethically problematic. For instance, the trolley problem dilemma was used to probe the differences between doing and allowing harm, with applications to bioethics, particularly abortion (Foot, 1967). Several philosophers describe being inspired by texts communicating empirical work, seeking to provide explanations for empirical observations (P1, P2, P16, P18) as well as subjecting the practices and products of the empirical sciences to critical inquiry (P2, P7, P12, P13, P18).
What do philosophers want out of their ideas? Once philosophers identify puzzles from conversations and texts, they aim to develop ideas which make progress on these puzzles. Progress is conceived of in many ways: “understand[ing] some part of the world better” (P3), working through new ways to think about problems (P17), and better understanding the current ways we think — for instance, by making implicit assumptions explicit and recognizing the implications of propositions (P7). Some philosophers describe a developed philosophical idea as a “picture” (P9, P10) which organizes subideas in a systematic way, allowing one to clearly see the main point(s). This often requires “conceptual engineering” (P6): challenging, disassembling, and rebuilding the ways in which we think.
@@
The role of conversation in philosophical development. Conversations with fellow philosophers are central to evaluating the coherence of ideas (P21), raising connections to other ideas and problems (P5), and collecting feedback (P3, P10). Conversations may force philosophers to explain and justify ideas they may have taken for granted (P1). Conversation helps philosophers gain confidence that their ideas are good intellectual contributions (P2, P21). Philosophers even simulate conversations in their head, taking on various positions for and against their ideas (P1, P12). Good philosophical conversation requires several conditions. The interlocutor should be charitable — genuinely listening to and working through ideas (P1, P12), and trusting (P6, P14) — but also willing to boldly push ideas forward (P3) and take intellectual risks (P18). Conversations may not be directed towards any clear goal; interlocutors must be able to “riff off each other” (P8) and be willing to operate without a preset agenda (P3, P4). This requires interlocutors to be curious about addressing problems (P21); it should be a collaborative effort, rather than a combative debate (P3, P7).